# Beyond boundaries: Gary Lorden’s groundbreaking contributions to sequential analysis

### v1

@@ 2.5 Lorden’s 2-SPRT and the Kiefer–Weiss Minimax Optimality
Suppose that based on a sequence of independent observations $\{X_{n}\}_{n\geqslant 1}$ with common parametric density $f_{\theta}$ one wishes to test the hypothesis $\operatorname{\mathcal{H}}_{0}:\theta=\theta_{0}$ versus $\operatorname{\mathcal{H}}_{1}:\theta=\theta_{1}$ ($\theta_{0}<\theta_{1}$) with error probabilities at most $\alpha_{0}$ and $\alpha_{1}$. Even though the SPRT has the remarkable optimality property of minimizing the expected sample size for both statistical hypotheses ${\mathsf{E}}_{\theta_{i}}[T]$, $i=0,1$, its performance may be poor when the true parameter value $\theta=\vartheta\in(\theta_{0},\theta_{1})$ differs from putative ones $\theta_{0}$ or $\theta_{1}$. Its expected sample size ${\mathsf{E}}_{\vartheta}[T]$ can be even much larger than that of the fixed sample size Neyman-Pearson test. See, e.g., Section 5.2 in Tartakovsky, Nikiforov, and Basseville (2015). Much work has been directed toward finding sequential tests that reduce the expected sample size of the SPRT for parameter values between the hypotheses.
@@
denote the expected sample size of an optimal test in the class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ in the worst-case scenario. The problem of finding a test $\delta_{0}=(T_{0},d_{0})$ such that $\sup_{\theta}{\mathsf{E}}_{\theta}[T_{0}]={\mathsf{ESS}}(\alpha_{0},\alpha_{1})$ subject to the error probability constraints $\alpha_{0}$ and $\alpha_{1}$ is known as the Kiefer–Weiss problem. No strictly optimal test has been found so far. Kiefer and Weiss (1957) presented structured results about tests which minimize the expected sample size ${\mathsf{E}}_{\theta}[T]$ at a selected point $\theta=\vartheta\in(\theta_{0},\theta_{1})$, which is referred to as the modified Kiefer–Weiss problem. Weiss (1962) proved that the Kiefer–Weiss problem reduces to the modified problem in symmetric cases for normal and binomial distributions. Lorden (1976) made a valuable contribution to the modified Kiefer–Weiss problem for two not necessarily parametric hypotheses $\operatorname{\mathcal{H}}_{i}:{\mathsf{P}}={\mathsf{P}}_{i}$, $i=0,1$, when the observations $X_{1},X_{2},\dots$ are i.i.d. and their true probability distribution ${\mathsf{G}}$ may be different from ${\mathsf{P}}_{0}$ and ${\mathsf{P}}_{1}$. Lorden (1976) introduced a simple binary combination of one-sided SPRTs that he called 2-SPRT and proved that this test is third-order asymptotically optimal. Later, Lorden (1980) proved theorems that characterize the basic structure of optimal sequential tests for the modified Kiefer–Weiss problem. His work has generated several works related to both the modified Kiefer–Weiss problem and the original Kiefer–Weiss problem of minimizing the maximal expected sample size; see, e.g., Huffman (1983), Dragalin and Novikov (1987), and Tartakovsky, Nikiforov, and Basseville (2015, Section 5.3).
@@
$p_{i}({\mathbf{X}}^{n})=\begin{cases}\prod_{t=1}^{n}f_{i}(X_{t})&\quad\text{for}~{}i=0,1,\dots,N,\\ \prod_{t=1}^{n}g(X_{t})&\quad\text{for}~{}i=N+1\end{cases}$
with respect to ${\mathsf{Q}}^{n}$, where $f_{i}(X_{t})$ and $g(X_{t})$, $t\geqslant 1$, are corresponding densities for the $t$-th observation.
For $n\in\mathbb{N}$ and $i=0,1,\dots,N$, define the LR and LLR processes
$\Lambda_{i}(n)=\frac{{\mathrm{d}}{\mathsf{G}}^{n}}{{\mathrm{d}}{\mathsf{P}}_{i}^{n}}({\mathbf{X}}^{n})=\prod_{t=1}^{n}\frac{g(X_{t})}{f_{i}(X_{t})},\quad\lambda_{i}(n)=\log\Lambda_{i}(n)=\sum_{t=1}^{n}\log\left[\frac{g(X_{t})}{f_{i}(X_{t})}\right],$
@@
For a matrix of positive thresholds $\mathbf{A}=(a_{ij})$, introduce the Markov times
$T_{i}=\inf\left\{t:\min_{\stackrel{{\scriptstyle 0\leqslant j\leqslant N}}{{j\neq i}}}[\lambda_{j}(t)-a_{ji}]\geqslant 0\right\}$ (32)
and define the test $\delta^{\star}=(T^{\star},d^{\star})$ as
$T^{\star}=\min_{0\leqslant i\leqslant N}T_{i},\qquad d^{\star}=i\quad\mbox{if}\quad T^{\star}=T_{i}.$ (33)
Note that $T_{i}$ is the time of accepting the hypothesis $H_{i}$, and also that it is a straightforward modification of the MSPRT (13)-(14). Indeed, the procedure (33) can be equivalently represented as:
$\text{Stop at the first $n\geqslant 1$ such that for some~{}$i$,}\quad\frac{{\mathrm{d}}{\mathsf{G}}^{n}}{{\mathrm{d}}{\mathsf{P}}_{j}^{n}}({\mathbf{X}}^{n})\geqslant e^{a_{ji}}\quad\text{for all~{}$j\neq i$},$ (34)
and accept $H_{i}$ for the unique $i$ that satisfies these inequalities. Comparing with the MSPRT (12) we see that the likelihood ratios between the hypotheses $H_{j}$ and $H_{i}$ are now replaced with the likelihood ratios between the measures ${\mathsf{G}}$ and ${\mathsf{P}}_{i}$. Hence, we will call it the modified accepting MSPRT.
Using Wald’s likelihood ratio identity, it is easy to show that the probabilities of errors $\alpha_{ij}(\delta^{\star})={\mathsf{P}}_{i}(d^{\star}=j)$ of the modified MSPRT satisfy the inequalities
$\alpha_{ij}(\delta^{\star})\leqslant e^{-a_{ij}}{\mathsf{G}}(d^{\star}=j)\quad\text{for}~{}i,j=0,1,\ldots,N,i\neq j.$
See Tartakovsky, Nikiforov, and Basseville (2015, Lemma 5.3.1 (page 230)). In the symmetric case where $a_{ij}=a$ and ${\mathsf{G}}(d^{\star}=j)=N^{-1}$,
$\alpha_{ij}(\delta^{\star})\leqslant N^{-1}e^{-a}.$
The case of two hypotheses $\operatorname{\mathcal{H}}_{0}$ and $\operatorname{\mathcal{H}}_{1}$ ($N=1$) considered by Lorden (1976) is of special interest. Here the modified MSPRT reduces to two parallel one-sided SPRTs,
@@
Its stopping time is $T^{\star}=\min(T_{0},T_{1})$ and the terminal decision is $d^{\star}=\arg\min_{i=0,1}T_{i}$. Lorden (1976) called this test the $2$-SPRT. If $a_{i}=\log(1/\alpha_{i})$, $i=0,1$, then $\alpha_{i}(\delta^{\star})={\mathsf{P}}_{i}(d^{\star}\neq i)\leqslant\alpha_{i}$, i.e., this test belongs to class ${\mathbb{C}}(\alpha_{0},\alpha_{1})=\{\delta:\alpha_{0}(\delta)\leqslant\alpha_{0},\alpha_{1}(\delta)\leqslant\alpha_{1}\}$. These upper bounds may be rather conservative. For example, in the symmetric case ${\mathsf{G}}(d^{\star}=1)={\mathsf{G}}(d^{\star}=0)=1/2$, we have $\alpha_{i}(\delta^{\star})\leqslant\alpha_{i}/2$.
Let ${\mathsf{E}}_{{\mathsf{G}}}$ denote expectation under ${\mathsf{G}}$ and let $I_{i}={\mathsf{E}}_{{\mathsf{G}}}[\lambda_{i}(1)]$, $i=0,1$, denote Kullback–Leibler information numbers. The following theorem proved by Lorden (1976) establishes third-order asymptotic optimality of Lorden’s 2-SPRT for small probabilities of errors $\alpha_{i}$. Its proof is based on Bayesian arguments. It follows from Theorem 1 in Lorden (1977a), which was proved a year later.
@@
Let the observations $\{X_{n}\}_{n\geqslant 1}$ be i.i.d. under ${\mathsf{G}},{\mathsf{P}}_{0}$, and ${\mathsf{P}}_{1}$. Assume that the Kullback-Leibler information numbers $I_{0}$ and $I_{1}$ are positive and, in addition, the second-moment conditions ${\mathsf{E}}_{{\mathsf{G}}}|\lambda_{i}(1)|^{2}<\infty$, $i=0,1$, hold. Let $\alpha^{\star}_{0}(a_{0},a_{1})$ and $\alpha^{\star}_{1}(a_{0},a_{1})$ denote the error probabilities of the 2-SPRT $\delta^{\star}(a_{0},a_{1})=(T^{\star}(a_{0},a_{1}),d^{\star}(a_{0},a_{1}))$. Let ${\mathsf{ESS}}(a_{0},a_{1})$ denote infimum of the expected sample size ${\mathsf{E}}_{{\mathsf{G}}}[T]$ over all tests with ${\mathsf{P}}_{0}(d=1)\leqslant\alpha_{0}^{\star}(a_{0},a_{1})$ and ${\mathsf{P}}_{1}(d=0)\leqslant\alpha_{1}^{\star}(a_{0},a_{1})$. Then
${\mathsf{ESS}}(a_{0},a_{1})={\mathsf{E}}_{{\mathsf{G}}}[T^{\star}(a_{0},a_{1})]+o(1)\quad\text{as}~{}\min(a_{0},a_{1})\to\infty,$ (36)
@@
The importance of this result cannot be overstated since the very simple test proposed by Lorden is almost optimal. At the same time, the optimal test can also be computed using Bellman’s backward induction algorithm since the optimal sequential test is truncated, i.e., has bounded maximal sample size, as shown by Kiefer and Weiss (1957). For one-parameter exponential families, the optimal bounds are curved in the $(S_{n},n)$ plane, where $S_{n}=\sum_{t=1}^{n}X_{i}$, and finding them typically requires substantial computation. In contrast, Lorden’s 2-SPRT approximates optimal curved boundaries with simple linear ones, so the continuation region is a triangle, as shown in Figure 1, below.
Lorden (1976) performed an extensive performance analysis for testing the mean $\theta$ of the Gaussian distribution $X_{n}\sim{\mathcal{N}}(\theta,1)$ with the hypotheses $\operatorname{\mathcal{H}}_{i}:\theta=\theta_{i}$, ($i=0,1$) in the symmetric case where $\alpha_{0}=\alpha_{1}$ and ${\mathsf{G}}\sim{\mathcal{N}}(\theta_{\star},1)$, $\theta_{\star}=(\theta_{0}+\theta_{1})/2$. The conclusion is that the $2$-SPRT performs very closely to the optimal test with its curved boundaries obtained using backward induction. The efficiency depends on the error probabilities, but it was over $99\%$ in all his performed experiments. Similar results were obtained by Huffman (1983) for the exponential example $f_{\theta}(x)=\theta e^{-\theta x}$, $x\geqslant 0$, $\theta>0$. Here the $2$-SPRT has efficiency over $98\%$ and almost always over $99\%$ for a broad range of error probabilities and parameter values.
The results of Theorem 4 can be extended to the multiple hypothesis case. Specifically, the modified matrix SPRT defined in (32)–(34) is also third-order asymptotically optimal as ${\alpha_{\rm max}}=\max_{1\leqslant i\leqslant N}\alpha_{i}\to 0$ in the class of tests
@@
Lorden’s results have led to further research related to the minimax Kiefer–Weiss problem of establishing near-optimal solutions for the least favorable intermediate distribution ${\mathsf{G}}$ in the single-parameter exponential family.
@@
We now consider the parametric case ${\mathsf{G}}={\mathsf{P}}_{\theta}$, ${\mathsf{P}}_{i}={\mathsf{P}}_{\theta_{i}}$ where the hypotheses are $H_{0}:\theta=\theta_{0}$ and $H_{1}:\theta=\theta_{1}$, $\theta_{0}<\theta_{1}$. Let $\theta$ be an arbitrary point belonging to the interval $(\theta_{0},\theta_{1})$ and let $\delta^{\star}(\theta)=(d^{\star}(\theta),T^{\star}(\theta))$ denote the $2$-SPRT tuned to $\theta$. In other words, $T^{\star}(\theta)=\min(T_{0}^{\theta},T_{1}^{\theta})$, where the $T_{i}^{\theta}$’s are defined by (35) with the LLRs $\lambda_{i}^{\theta}(n)=\log[{\mathrm{d}}{\mathsf{P}}_{\theta}^{n}/{\mathrm{d}}{\mathsf{P}}_{\theta_{i}}^{n}]({\mathbf{X}}^{n})$, $i=0,1$, tuned to $\theta$.
@@
Theorem 4 implies that the $2$-SPRT $\delta^{\star}(\theta)$ is third-order asymptotically optimal in terms of minimizing ${\mathsf{E}}_{\theta}[T]$ at the intermediate point $\theta\in(\theta_{0},\theta_{1})$ when the second moments ${\mathsf{E}}_{\theta}|\lambda_{0}^{\theta}(1)|^{2}$ and ${\mathsf{E}}_{\theta}|\lambda_{1}^{\theta}(1)|^{2}$ are finite and the thresholds $a_{i}$ are selected in such a way that the error probabilities are either exactly equal to the given numbers $\alpha_{i}$ or at least close to $\alpha_{i}$, the latter requirement being a difficult task. However, setting $a_{i}=|\log\alpha_{i}|$ embeds the $2$-SPRT into the class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ and Theorem 4 suggests that if one can find a nearly least favorable point $\theta^{*}$, i.e. $\theta^{*}$ can be selected so that $\sup_{\theta}{\mathsf{E}}_{\theta}[T^{\star}(\theta^{*})]\approx{\mathsf{E}}_{\theta^{*}}[T^{\star}(\theta^{*})]$, then $\delta^{\star}(\theta^{*})$ is an approximate solution to the Kiefer–Weiss problem of minimizing $\sup_{\theta}{\mathsf{E}}_{\theta}[T]$.
@@
Consider the single-parameter exponential family $\{f_{\theta}(x),\theta\in\Theta\}$ assuming that the common density $f_{\theta}(x)$ of the i.i.d. observations $X_{1},X_{2},\dots$ is such that
$\frac{f_{\theta}(X_{n})}{f_{\widetilde{\theta}}(X_{n})}=\exp\left\{(\theta-\widetilde{\theta})X_{n}-(b(\theta)-b(\widetilde{\theta}))\right\},\quad n=1,2,\dots$ (37)
where $b(\theta)$ is a convex and infinitely differentiable function on $\widetilde{\Theta}\subset\Theta$. A simple calculation shows that ${\mathsf{E}}_{\theta}[X_{1}]=\overset{\bm{.}}{b}(\theta)$ , $\sigma^{2}_{\theta}={\mathsf{var}}_{\theta}(X_{1})=\ddot{b}(\theta)$, and the Kullback–Leibler numbers are
$I(\theta,\theta_{i})=(\theta-\theta_{i})\overset{\bm{.}}{b}(\theta)-(b(\theta)-b(\theta_{i})).$
Without loss of generality we assume that $\inf_{\theta\in\widetilde{\Theta}}\sigma^{2}_{\theta}>0$, $\sup_{\theta\in\widetilde{\Theta}}\sigma^{2}_{\theta}<\infty$.

### v2

@@ 2.3 Lorden’s 2-SPRT and the Kiefer–Weiss Minimax Optimality
Suppose that based on a sequence of independent observations $\{X_{n}\}_{n\geqslant 1}$ with common parametric density $f_{\theta}$ one wishes to test the hypothesis $\operatorname{\mathcal{H}}_{0}:\theta=\theta_{0}$ versus $\operatorname{\mathcal{H}}_{1}:\theta=\theta_{1}$ ($\theta_{0}<\theta_{1}$) with error probabilities at most $\alpha_{0}$ and $\alpha_{1}$. Even though the SPRT has the remarkable optimality property of minimizing the expected sample size for both statistical hypotheses ${\mathsf{E}}_{\theta_{i}}[T]$, $i=0,1$, its performance may be poor when the true parameter value $\theta=\vartheta\in(\theta_{0},\theta_{1})$ differs from putative ones $\theta_{0}$ or $\theta_{1}$. Its expected sample size ${\mathsf{E}}_{\vartheta}[T]$ can be even much larger than that of the fixed sample size of the Neyman-Pearson test. See, e.g., Section 5.2 in Tartakovsky, Nikiforov, and Basseville (2015). Much work has been directed toward finding sequential tests that reduce the expected sample size of the SPRT for parameter values between the hypotheses.
@@
denote the expected sample size of an optimal test in the class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$ in the worst-case scenario. The problem of finding a test $\delta_{0}=(T_{0},d_{0})$ such that $\sup_{\theta}{\mathsf{E}}_{\theta}[T_{0}]={\mathsf{ESS}}(\alpha_{0},\alpha_{1})$ subject to the error probability constraints $\alpha_{0}$ and $\alpha_{1}$ is known as the Kiefer–Weiss problem. No strictly optimal test has been found so far. Kiefer and Weiss (1957) presented structured results about tests which minimize the expected sample size ${\mathsf{E}}_{\theta}[T]$ at a selected point $\theta=\vartheta\in(\theta_{0},\theta_{1})$, which is referred to as the modified Kiefer–Weiss problem. Weiss (1962) proved that the Kiefer–Weiss problem reduces to the modified problem in symmetric cases for normal and binomial distributions. Lorden (1976) made a valuable contribution to the modified Kiefer–Weiss problem for two not necessarily parametric hypotheses $\operatorname{\mathcal{H}}_{i}:{\mathsf{P}}={\mathsf{P}}_{i}$, $i=0,1$, when the observations $X_{1},X_{2},\dots$ are i.i.d. and their true distribution ${\mathsf{P}}_{2}$ may be different from ${\mathsf{P}}_{0}$ and ${\mathsf{P}}_{1}$. Lorden (1976) introduced a simple combination of one-sided SPRTs, called the 2-SPRT, and proved that it is third-order asymptotically optimal. Later, Lorden (1980) proved theorems that characterize the basic structure of optimal sequential tests for the modified Kiefer–Weiss problem. His work has generated several works related to both the modified Kiefer–Weiss problem and the original Kiefer–Weiss problem of minimizing the maximal expected sample size; see, e.g., Huffman (1983), Dragalin and Novikov (1987), and Tartakovsky, Nikiforov, and Basseville (2015, Section 5.3).
@@
For $n\in\mathbb{N}$ and $i=0,1$, define the LR and LLR processes
$\Lambda_{i}(n)=\frac{{\mathrm{d}}{\mathsf{P}}_{2}^{n}}{{\mathrm{d}}{\mathsf{P}}_{i}^{n}}({\mathbf{X}}^{n})=\prod_{t=1}^{n}\frac{f_{2}(X_{t})}{f_{i}(X_{t})},\quad\lambda_{i}(n)=\log\Lambda_{i}(n)=\sum_{t=1}^{n}\log\left[\frac{f_{2}(X_{t})}{f_{i}(X_{t})}\right],$
@@
Define two parallel one-sided SPRTs
@@
The stopping time of Lorden’s 2-SPRT (Lorden 1976) is $T^{\star}=\min(T_{0},T_{1})$ and the terminal decision is $d^{\star}=\arg\min_{i=0,1}T_{i}$. If $a_{i}=\log(1/\alpha_{i})$, $i=0,1$, then $\alpha_{i}(\delta^{\star})={\mathsf{P}}_{i}(d^{\star}\neq i)\leqslant\alpha_{i}$, i.e., this test belongs to class ${\mathbb{C}}(\alpha_{0},\alpha_{1})=\{\delta:\alpha_{0}(\delta)\leqslant\alpha_{0},\alpha_{1}(\delta)\leqslant\alpha_{1}\}$. These upper bounds may be rather conservative. For example, in the symmetric case ${\mathsf{P}}_{2}(d^{\star}=1)={\mathsf{P}}_{2}(d^{\star}=0)=1/2$, we have $\alpha_{i}(\delta^{\star})\leqslant\alpha_{i}/2$.
Let ${\mathsf{E}}_{2}$ denote expectation under ${\mathsf{P}}_{2}$, and let $I_{i}={\mathsf{E}}_{2}[\lambda_{i}(1)]$, $i=0,1$, denote Kullback–Leibler information numbers. The following theorem, proved by Lorden (1976), establishes third-order asymptotic optimality of Lorden’s 2-SPRT for small probabilities of errors $\alpha_{i}$. Its proof is based on Bayesian arguments. This theorem emerges from Theorem 1 in Lorden (1977a), which was proved a year later.
@@
Let the observations $\{X_{n}\}_{n\geqslant 1}$ be i.i.d. under ${\mathsf{P}}_{i}$, $i=0,1,2$. Assume that the Kullback-Leibler information numbers $I_{0}$ and $I_{1}$ are positive and, in addition, the second-moment conditions ${\mathsf{E}}_{2}|\lambda_{i}(1)|^{2}<\infty$, $i=0,1$, hold. Let $\alpha^{\star}_{0}(a_{0},a_{1})$ and $\alpha^{\star}_{1}(a_{0},a_{1})$ denote the error probabilities of the 2-SPRT $\delta^{\star}(a_{0},a_{1})=(T^{\star}(a_{0},a_{1}),d^{\star}(a_{0},a_{1}))$. Let ${\mathsf{ESS}}(a_{0},a_{1})$ denote infimum of the expected sample size ${\mathsf{E}}_{2}[T]$ over all tests with ${\mathsf{P}}_{0}(d=1)\leqslant\alpha_{0}^{\star}(a_{0},a_{1})$ and ${\mathsf{P}}_{1}(d=0)\leqslant\alpha_{1}^{\star}(a_{0},a_{1})$. Then
${\mathsf{ESS}}(a_{0},a_{1})={\mathsf{E}}_{2}[T^{\star}(a_{0},a_{1})]+o(1)\quad\text{as}~{}\min(a_{0},a_{1})\to\infty,$ (29)
@@
The significance of this result cannot be overstated, as Lorden’s simple test is nearly optimal. Simultaneously, the optimal test can be computed using Bellman’s backward induction algorithm since the optimal sequential test is truncated, meaning it has a bounded maximal sample size, as demonstrated by Kiefer and Weiss (1957). For one-parameter exponential families $\{{\mathsf{P}}_{\theta},\theta\in\theta\}$, the optimal bounds exhibit curvature in the $(S_{n},n)$ plane, where $S_{n}=\sum_{t=1}^{n}X_{i}$, and determining them typically entails substantial computation. In contrast, Lorden’s 2-SPRT approximates optimal curved boundaries with simple linear ones, resulting in a continuation region shaped like a triangle, as illustrated in Figure 2.
Figure 2: The boundaries $h_{1}^{\theta}(n)$ and $h_{0}^{\theta}(n)$ of the $2$-SPRT (solid) and optimal boundaries (dashed) as functions of $n$. $S_{n}^{\theta}=S_{n}-n{\mathsf{E}}_{\theta}[X_{1}]$.
Lorden (1976) performed an extensive performance analysis for testing the mean $\theta$ of the Gaussian distribution $X_{n}\sim{\mathcal{N}}(\theta,1)$ with the hypotheses $\operatorname{\mathcal{H}}_{i}:\theta=\theta_{i}$, ($i=0,1$) in the symmetric case where $\alpha_{0}=\alpha_{1}$ and ${\mathsf{P}}_{2}\sim{\mathcal{N}}(\theta_{\star},1)$, $\theta_{\star}=(\theta_{0}+\theta_{1})/2$. The conclusion is that the $2$-SPRT performs very closely to the optimal test with its curved boundaries obtained using backward induction. The efficiency depends on the error probabilities, but it was over $99\%$ in all his performed experiments. Similar results were obtained by Huffman (1983) for the exponential example $f_{\theta}(x)=\theta e^{-\theta x}$, $x\geqslant 0$, $\theta>0$. Here the $2$-SPRT has efficiency over $98\%$ and almost always over $99\%$ for a broad range of error probabilities and parameter values.
The results of Theorem 4 can be extended to the multiple hypothesis case with $N+1$ hypotheses, $N>1$. Specifically, the modified matrix SPRT is also third-order asymptotically optimal as ${\alpha_{\rm max}}=\max_{0\leqslant i\leqslant N}\alpha_{i}\to 0$ in the class of tests
@@
Lorden’s results have sparked further research into the minimax Kiefer–Weiss problem, which aims to establish near-optimal solutions for the least favorable intermediate distribution ${\mathsf{P}}_{2}$ within the single-parameter exponential family. Consider the parametric case ${\mathsf{P}}_{2}={\mathsf{P}}_{\theta}$, ${\mathsf{P}}_{i}={\mathsf{P}}_{\theta_{i}}$ where the hypotheses are $\operatorname{\mathcal{H}}_{0}:{\mathsf{P}}_{0}={\mathsf{P}}_{\theta_{0}}$ and $\operatorname{\mathcal{H}}_{1}:{\mathsf{P}}_{1}={\mathsf{P}}_{\theta_{1}}$, $\theta_{0}<\theta_{1}$. Let $\theta$ be an arbitrary point belonging to the interval $(\theta_{0},\theta_{1})$ and let $\delta^{\star}(\theta)=(d^{\star}(\theta),T^{\star}(\theta))$ denote the $2$-SPRT tuned to $\theta$. In other words, $T^{\star}(\theta)=\min(T_{0}^{\theta},T_{1}^{\theta})$, where the $T_{i}^{\theta}$’s are defined by (28) with the LLRs $\lambda_{i}^{\theta}(n)=\log[{\mathrm{d}}{\mathsf{P}}_{\theta}^{n}/{\mathrm{d}}{\mathsf{P}}_{\theta_{i}}^{n}]({\mathbf{X}}^{n})$, $i=0,1$, tuned to $\theta$.
@@
Theorem 4 implies that the $2$-SPRT $\delta^{\star}(\theta)$ is third-order asymptotically optimal for minimizing ${\mathsf{E}}_{\theta}[T]$ at the intermediate point $\theta\in(\theta_{0},\theta_{1})$ when the second moments ${\mathsf{E}}_{\theta}|\lambda_{0}^{\theta}(1)|^{2}$ and ${\mathsf{E}}_{\theta}|\lambda_{1}^{\theta}(1)|^{2}$ are finite, and the thresholds $a_{i}$ are selected so that the error probabilities are either exactly equal, or at least close, to the given numbers $\alpha_{i}$, a challenging task. However, setting $a_{i}=|\log\alpha_{i}|$ embeds the $2$-SPRT into the class ${\mathbb{C}}(\alpha_{0},\alpha_{1})$, and Theorem 4 suggests that if one can find a nearly least favorable point $\theta^{*}$, i.e., $\theta^{*}$ can be selected so that $\sup_{\theta}{\mathsf{E}}_{\theta}[T^{\star}(\theta^{*})]\approx{\mathsf{E}}_{\theta^{*}}[T^{\star}(\theta^{*})]$, then $\delta^{\star}(\theta^{*})$ is an approximate solution to the Kiefer–Weiss problem of minimizing $\sup_{\theta}{\mathsf{E}}_{\theta}[T]$.
@@
For the single-parameter exponential family with density $\{f_{\theta}(x),\theta\in\Theta\}$, where
$\frac{f_{\theta}(x)}{f_{\tilde{\theta}}(x)}=\exp\left\{(\theta-\tilde{\theta})x-(b(\theta)-b(\tilde{\theta}))\right\},$ (30)